# GPU-based track-finding for the J-PARC muon g-2/EDM experiment

### v1

@@ 7 Analysis
Using the simulator and the track finding tool, the track finding efficiency in a multiple-track condition is estimated. The efficiency, $\varepsilon$, is defined as follows:
@@
The generated momentum distribution and the reconstructed momentum distribution for different pileup rates of the positron tracks are shown in Figure 6 respectively.
Figure 6: (a) Generated momentum distribution (b) Momentum distribution of found positron tracks
Figure 7 shows the track-finding efficiency for a single track as a function of the positron momentum under different track rate conditions. Even under the highest track rate condition of 6 tracks/ns, the finding efficiency remains above 90% in the momentum range of $200~\text{MeV}/c<p<275~\text{MeV}/c$.
Figure 7: Momentum dependence of track finding efficiency with different conditions of track rates from 0.06 tracks/ns to 6 tracks/ns.
Tracking detectors can occasionally reconstruct false hits and tracks, known as ghost hits and ghost tracks. To evaluate their impact, the contribution of ghost hits in reconstructed tracks is studied for single-track events. It is found that approximately 50% of the tracks are found without any ghost hits, and the majority of the remaining tracks contain less than 20% ghost hit contribution. Some of the contributions from ghost hits are removed during the track-fitting process, and with a more optimized track-finding algorithm, they can be further reduced.
Figure 8: The fraction of ghost hits in the found tracks.
The computation times for GPU activities, including kernel execution and data transfers between host and device (both Host-to-Device and Device-to-Host), are summarized in Table 1. These measurements were obtained using the nvprof profiling tool for track rates ranging from 0.06 to 6 tracks/ns on a Tesla P100-PCIE GPU with 16 GiB of memory, using $10^{4}$ simulated muon decay events.
@@
Table 1: Approximate Computation time for GPU kernel execution, CPU, and memory transfers at varying muon decay rates of $10^{4}$ muons, measured using nvprof [32].
@@
8 Summary and conclusion
Muon g-2/EDM experiment at J-PARC measures the $a_{\mu}$ and EDM with a new approach to reach the uncertainty of 460 ppb. One of the crucial part of the analysis lies in the reconstruction of the positron tracks. These positron tracks are reconstructed from the hits in silicon strip detector. In the experiment, data collection (acquisition) can run the expected speed of $10^{4}$ muons every 7 minutes(24 muons/sec/CPU). With 1000 CPUs, the data processing rate reaches about $10^{5}$ muons per second. The software must handle data from approximately $10^{6}$ muon decays per second (see Section 5).
With the implementation of the GPU for specification algorithm in track finding, we are able to accelerate the computation time by a factor of around 7 $\times$ for the pileup rate of 0.06 muon decay/ns. (see Table 1). The efficiency is for the three pileup rate ( 0.06 tracks/ns, 0.6 tracks/ns and, 6 tracks/ns) is 90 - 95 %. We evaluated the ghost hit contamination at the lowest pile-up rate. As shown in Fig. 8, approximately 50% of the tracks are reconstructed without any ghost hits. Among the remaining tracks, the majority exhibit a ghost hit contribution of less than 20%. As discussed in Section 7, Figure 9 shows that GPU-based processing significantly speeds up the track finding process. This improvement is attributed to the GPU’s higher compute capability, greater memory bandwidth, and large number of parallel processing cores, offering a clear advantage over traditional CPU-based methods. In particular, the NVIDIA RTX 4090 and H200 achieves over a tenfold improvement in performance compared to CPU-based methods.
In the simulation framework, track fitting is one of the most time-consuming steps. It uses input from either RecoHits or StripClusters, which are processed signals from the tracking detectors. The track parameters are estimated at the position of the first detector hit using the Kalman filter algorithm, as implemented in the Genfit2 toolkit [27]. To further speed up the simulation, the next step in our study could be to implement the track fitting algorithm on the GPU. By rewriting both the track finding and track fitting algorithms to run entirely on the GPU, we can significantly accelerate the overall simulation process.

### v2

@@ 7 Analysis
Using the simulator and the track-finding tool, the track-finding efficiency in a multiple-track condition is estimated. The efficiency, $\varepsilon$, is defined as follows:
@@
The generated momentum distribution and the reconstructed momentum distribution for different pileup rates of the positron tracks are shown in figure 5 respectively. The plot is marked by lower efficiencies below about 80 MeV/c due to tracks without sufficient hits for reconstruction and a significant dip at about 150 Mev/c due to tracks that are disjoint as they cross the void central volume of the detector (refer figure 1).
Figure 5: (a) Generator level momentum distribution of decay positrons from Geant4 (b) The momentum distribution from decay positron tracks identified by the track-finding algorithm.
Figure 6 shows the track-finding efficiency for a single track as a function of the positron momentum under different track rate conditions. At the currently proposed intensity of $10^{4}$ muons/pulse, equivalent to $2\times 10^{5}$ muons per second leading to an estimated highest decay rate of 1.5 decays/ns. The study shows that even for a much higher track rate of 6 tracks/ns, the finding efficiency still remains above 90% in the momentum range of $200~\text{MeV}/c<p<275~\text{MeV}/c$, the window that is also useful for muon g-2 analysis.
Figure 6: Momentum dependence of track-finding efficiency with different conditions of track rates from 0.06 tracks/ns to 6 tracks/ns.
Tracking detectors can occasionally reconstruct false hits and tracks, known as ghost hits. Ghost hits are incorrect combinations of R-sensors hits and Z-sensors hits used in track reconstruction which can impact the performance of the track reconstruction. To evaluate their impact, the contribution of ghost hits in reconstructed tracks is studied for single-track events. It is found that approximately 50% of the tracks are found without any ghost hits, and the majority of the remaining tracks contain less than 20% ghost hit contribution. Some of the contributions from ghost hits are removed during the track-fitting process, and with a more optimized track-finding algorithm, they can be further reduced.
Figure 7: The fraction of ghost hits identified by the track-finding algorithm.
The computation times for GPU activities, including kernel execution and data transfers between host and device (both Host-to-Device and Device-to-Host), are summarized in table 1. These measurements were obtained using the nvprof profiling tool for track rates ranging from 0.06 to 6 tracks/ns on a Tesla P100-PCIE GPU with 16 GiB of memory, using $10^{4}$ simulated muon decay events  [25].
@@
Table 1: Approximate computation time for GPU kernel execution, CPU, and memory transfer rates between host to device and device to host, at varying muon decay rates of $10^{4}$ muons, measured using nvprof, a GPU performance profiling application  [25].
@@
It is clear at this point that the implementation of the track-finding on a GPU can lead to a speedup improvement between 7-11 $\times$ compared to the CPU version depending on the GPU architecture.
8 Summary and conclusions
[…]
With the implementation of a GPU-based algorithm for track-finding, we are able to accelerate the computation time by a factor of around 7 - 11 $\times$ for the pileup rate of 0.06 muon decay/ns (see table 1), depending on the architecture. This is a forward-looking development expecting that GPUs will be much more commonly used in near future. The efficiency for different pileup rates (0.06 tracks/ns, 0.6 tracks/ns and, 6 tracks/ns) is estimated to be 90-95%, matching well with the expectations from the CPU-based algorithm. We evaluated the ghost hit contamination at the lowest pile-up rate. As shown in figure 7, approximately 50% of the tracks are reconstructed without any ghost hits. Among the remaining tracks, the majority exhibit a ghost hit contribution of less than 20%. As discussed in section 7 and as inferred from figure 8, GPU-based processing significantly speeds up the track-finding process.
This improvement is attributed to the GPU’s higher compute capability, greater memory bandwidth, and large number of parallel processing cores, offering a clear advantage over traditional CPU-based methods. In particular, the H200 achieves about 11 $\times$ improvement in performance compared to CPU-based methods.
In the simulation framework, track fitting is one of the most time-consuming steps. It uses input from either RecoHits or StripClusters, which are processed signals from the tracking detectors. The track parameters are estimated at the position of the first detector hit using the Kalman filter algorithm, as implemented in the Genfit2 toolkit [11]. To further speed up the simulation, the next step in our study could be to implement the track fitting algorithm on the GPU. By integrating both the track-finding and track-fitting algorithms to run entirely on the GPU, we expect to significantly accelerate the overall simulation process.